OpenStudy (anonymous):

-12 ( -4x + 6 ) -x + 7 - (-9x) Show your work please.

OpenStudy (anonymous):

how about i guide you through, sound good? :]

OpenStudy (anonymous):

Yep sounds good to me.

OpenStudy (anonymous):

ok so the distributive property with -12(-4x +6x) you get 48 -72

OpenStudy (anonymous):

-x +7 +9x because -1( -9x ) turns that 9 into a positive so 48 - 72 - x + 7 + 9x combine the like terms you get?

OpenStudy (anonymous):

-12 (-4x + 6) = -12 . -4x + -12 . 6 = 48x -72 - (-9x) = 9x 48x - x + 9x - 72 = 56x - 72 = 7x - 9

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

idk where your getting all these equal signs from? i got 48 - 72 - x + 7 + 9x -17 +8x but you never had that equal to anything

OpenStudy (anonymous):

eh.., i'm forget.., it should be 48x - x + 9x - 72 + 7 = 56x - 65 about i get 48x is because -12 times -4x is 48x

OpenStudy (anonymous):

Woah, so confused guys.

OpenStudy (anonymous):

48x - 72 -x + 7 + 9 = 0 combine the like terms

OpenStudy (anonymous):

so, we would use the " eye brow method here. -12 x -4 = ? and -12 x 6 = ? correct?

OpenStudy (anonymous):

so it would be, 48x - 72 -1x + 7 + 9 = 0 ? Then from there we would add the varibles or seperate terms?

OpenStudy (anonymous):

combining like terms, yes

OpenStudy (anonymous):

like... 48x - 1x = 47x -72 + 7 + 9 = 56 ?

OpenStudy (anonymous):

- 56 **

OpenStudy (anonymous):

so it would end up being.... 47x - 56 ?

OpenStudy (anonymous):

yeah and i'm guessing you need to solve for x?

OpenStudy (anonymous):

No, um. Simplify I guess you could call it. Its just add em up. write it. We havent got that far yet. Thanks though. I have a better understanding now :D

OpenStudy (anonymous):

i want know how you get 9.., i think -(-9x) is - times -9x = 9x ( - x - = + ), and where is the "x" variable?? O.O

OpenStudy (anonymous):

All varibles without a stated number are always 1.